OpenStudy (anonymous):

Integral e^x/(1+e^2x) dx

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

arctg(e^x) +C

OpenStudy (shayaan_mustafa):

does its integral is defined?

OpenStudy (anonymous):

Integral value = arctg(e^x) +C

OpenStudy (anonymous):

simple substitution e^x = t

OpenStudy (anonymous):

\[\int\limits dt/1+t ^{2} = arctg (t) +C\]